# Neat: Nonlinear Parameter-efficient Adaptation of Pre-trained Models

### v2

@@ 6 Conclusion
In this work, we propose Neat, a novel parameter-efficient fine-tuning (PEFT) method that introduces nonlinear transformations to enhance model adaptation while maintaining efficiency. By incorporating a lightweight neural network that models cumulative weight updates as functions of the pre-trained weights, Neat effectively captures complex, nonlinear structures in the weight space, allowing for more expressive and accurate adaptation to downstream tasks. Our theoretical analysis supports the efficacy of Neat, demonstrating that it can achieve greater or equivalent expressiveness compared to existing LoRA, a popular and state-of-the-art PEFT method, with fewer number of parameters. Through extensive experiments on four benchmarks encompassing over twenty datasets with various pre-trained backbones, Neat demonstrated superior performance on both NLP and vision tasks compared to existing state-of-the-art methods.
Appendix
Appendix A Details of Theoretical Results
@@
Take $\mathbf{A}^{\#}=\sigma(\mathbf{W}^{0}\bm{\Theta}_{1}^{*})\in\mathbb{R}^{d_{1}\times r}$ and $\mathbf{B}^{\#}=\bm{\Theta}_{2}^{*}\in\mathbb{R}^{r\times d_{2}}$, where
$(\bm{\Theta}_{1}^{*},\bm{\Theta}_{2}^{*})=\operatorname*{arg\,min}_{\bm{\Theta}_{1}\in\mathbb{R}^{d_{2}\times r},\bm{\Theta}_{2}\in\mathbb{R}^{r\times d_{1}}}\mathcal{L}(\mathcal{D}_{\text{train}};\mathbf{W}^{0}+f(\mathbf{W}^{0};(\bm{\Theta}_{1},\bm{\Theta}_{2}))).$
The conclusion follows from
@@
A.2 Theoretical Analysis of Neat under sinusoid activation function
Here we consider a sinusoid activation function $\sigma_{\textnormal{p}}(x)=\sin(2\pi x)$ [Gashler and Ashmore, 2014] and design $f(\mathbf{W}^{0};\bm{\theta})=\sigma_{\textnormal{p}}(\mathbf{W}^{0}\bm{\Theta}_{1})\bm{\Theta}_{2}$ with $\bm{\theta}=(\bm{\Theta}_{1},\bm{\Theta}_{2})$. With this periodic activation function, we can show a stronger result that Neat has expressivity (almost) greater than or equal to a LoRA with more parameters when $d_{1}\gg d_{2}$.
Proposition A.1 (Expressivity of Neat with Sine Activation).
@@ A.2 Theoretical Analysis of Neat under sinusoid activation function
Proposition A.1 shows that the class of updates $\Delta\mathbf{W}=\sigma_{\textnormal{p}}(\mathbf{W}^{0}\bm{\Theta}_{1})\bm{\Theta}_{2}$ by Neat with $2rd_{2}$ parameters is dense in the class of updates $\Delta\mathbf{W}=\mathbf{A}\mathbf{B}$ by LoRA with $r(d_{1}+d_{2})$ parameters. When $d_{2}\ll d_{1}$, this shows better parameter efficiency of Neat.
Examining the proof of Proposition A.1, it is straightforward to show that the result holds for any continuous and periodic activation function whose range contains an open interval centered at 0.
@@
This proof relies on Kronecker’s theorem (Theorem 7.9 in Apostol [1990]) from number theory, which shows that for all $j\in\mathbb{R}^{q}$, the fractional parts of $(ct_{1},ct_{2},\dots,ct_{q})^{\top}$ is dense in $[0,1]^{q}$ over $c\in\mathbb{R}$, as long as $t_{1},\dots,t_{q}$ are linearly independent over the rationals.
@@
|$\absolutevalue{\{c_{j}\mathbf{W}_{j^{*}}\}-\frac{\arcsin(\mathbf{A}_{j})}{2\pi}}\leq\epsilon^{\prime},$
@@
Appendix B Additional Related Work
B.1 Additive PEFT Methods
Additive PEFT methods [Chronopoulou et al., 2023, Edalati et al., 2022, Lester et al., 2021, Wang et al., 2024a, Liu et al., 2022] introduces a small set of additional trainable parameters strategically placed within the model. One of the most prominent additive PEFT approaches is Adapter [Chronopoulou et al., 2023, Edalati et al., 2022, Zhao et al., 2022], which involves inserting small adapter layers between pre-trained weight blocks. Prompt Tuning [Wang et al., 2024a, Lester et al., 2021, Vu et al., 2021, Li and Liang, 2021] is another technique, where learnable vectors, or ”soft prompts,” are prepended to the input sequence without modifying the model’s weights. This method is particularly effective for large-scale models and has inspired variants such as Prefix Tuning [Li and Liang, 2021].
B.2 Selective PEFT Methods
Selective PEFT focuses on optimizing the fine-tuning process by selectively adjusting a subset of the model’s parameters rather than introducing additional ones. For instance, Diff Pruning [Guo et al., 2020] uses a learnable binary mask to select parameters for fine-tuning. Similarly, FishMask [Sung et al., 2021] and Fish-Dip [Das et al., 2023] leverage Fisher information to determine parameter importance and identify the most crucial ones for updates. Additionally, BitFit [Zaken et al., 2021] fine-tunes only the bias terms in the model, significantly reducing the number of trainable parameters.
B.3 Hybrid PEFT method
Hybrid PEFT methods aim to combine the strengths of various existing PEFT techniques to enhance model performance across diverse tasks. UniPELT [Mao et al., 2021] integrates LoRA, prefix-tuning, and adapters within each Transformer block, employing a gating mechanism to determine which module should be active during fine-tuning. S4 [Chen et al., 2023] further explores the design space by partitioning layers into groups and assigning different PEFT methods to each group. Additionally, NOAH [Zhang et al., 2022] and AUTOPEFT [Zhou et al., 2024] leverage neural architecture search (NAS) to automatically discover optimal combinations of PEFT techniques tailored to specific tasks.
Appendix C Baselines
Our baselines are constructed on a task basis. Specifically, for each task, the proposed Neat is compared with representative baselines from the corresponding domain, as listed below.
  • For both Commonsense Reasoning and Arithmetic Understanding, following Wang et al. [2024c], LoRA [Hu et al., 2021a], PiSSA [Meng et al., 2024] and MiLoRA [Wang et al., 2024c] are employed as baselines. Neat is applied to query, key, value, MLP up and MLP down layers.
  • For Natural Language Understanding, we follow the setup from prior works [Gao et al., 2024, Wu et al., 2024a] that evaluate various representative PEFT methods, including LoRA [Hu et al., 2021a], Adapter Houlsby et al. [2019], BitFit [Zaken et al., 2021], RED [Wu et al., 2024b], DoRA [Liu et al., 2024], ReFT Wu et al. [2024a], and FourierFT [Gao et al., 2024].
  • For Image Classification, we follow the setting of Gao et al. [2024] and take linear probing (LP), LoRA [Hu et al., 2021a] and FourierFT [Gao et al., 2024] as baselines. Neat is applied to the query and value layers.
Appendix D Hyperparameters
We provide the specific hyperparameters used in our experiments to ensure reproducibility. For most of our experiments, we use the standard implementation of Neat, which we refer to as vanilla Neat. The neural network architecture in vanilla Neat consists of only two layers: an input layer and an output layer. We selecte this approach because vanilla Neat offers the benefits of simplicity in implementation, a low parameter count, and sufficient adaptation power. Nonetheless, we dedicate Section 4.4 and Appendix F to exploring more complex adaptation networks and their effect on performance.
D.1 Image Classification
Hyperparameters for Neat for Fig. 4 are provided in Table 6. We tune the classification head and the backbone separately and provide detailed settings for each dataset. All weight decay values are not tuned and follow the settings from Gao et al. [2024]. The scaling factor $s$ is set to $1.0$. The hidden layer dimension $r$ for MHSA is set to 7 in the QV-setting, while both hidden layer dimensions for MHSA and MLP are set to 2 in the QV-MLP-setting described in Section 5.
Table 6: Hyperparameter of image classification for Neat.
@@ D.1 Image Classification
Learning Rate (Neat) 5E-3 1E-2 5E-3 1E-2 5E-3 1E-2 5E-3 5E-3
@@
Learning Rate (Neat) 5E-3 5E-3 5E-3 1E-2 5E-3 5E-3 1E-2 5E-3
@@
D.2 Natural Language Understanding
We provide used hyper-parameters for Neat in natural language understanding on the GLUE benchmark in Table 7 and Table 8. The reported results are obtained when using a depth of 6 for Neat. The learning rates for the head and the backbone are tuned separately. The scaling factor $s$ is searched in $\{0.01,0.1,1.0\}$. For reproducibility, we fix the seed as 0. The hidden layer dimension $r$ is set to 8 in Neat-L and 1 in Neat-S. More specifically, we apply Neat to all layers in RoBERTa-base for Neat-L, while only applying Neat to layers $\{4,5,6,7,8,9,10,11\}$ for Neat-S to reduce the number of trainable parameters. The seed is fixed for reproducibility.
Table 7: Hyperparameter of GLUE benchmark for Neat-L.
Hyperparameter STS-B RTE MRPC CoLA SST-2 QNLI MNLI QQP
Optimizer AdamW
LR Schedule Linear
Learning Rate (Neat) 5E-3 5E-3 5E-3 1E-3 5E-3 1E-3 5E-3 5E-3
Learning Rate (Head) 5E-3 5E-3 5E-3 1E-3 5E-3 1E-3 5E-3 5E-3
Scaling 0.1 0.01 0.01 0.1 0.01 0.01 0.01 0.01
Max Seq. Len 512 512 512 512 512 512 512 512
Batch Size 64 32 64 64 32 32 32 64
Table 8: Hyperparameter of GLUE benchmark for Neat-S.
Hyperparameter STS-B RTE MRPC CoLA SST-2 QNLI MNLI QQP
Optimizer AdamW
LR Schedule Linear
Learning Rate (Neat) 5E-3 1E-3 5E-3 5E-3 5E-3 1E-3 5E-3 1E-3
Learning Rate (Head) 1E-3 1E-3 5E-3 1E-3 5E-3 1E-3 5E-3 1E-3
Scaling 0.1 1.0 0.01 0.1 0.01 0.1 0.01 1.0
Max Seq. Len 512 512 512 512 512 512 512 512
Batch Size 64 32 64 64 32 32 32 64
D.3 Commonsense Reasoning
Table 9: Hyperparameter of commonsense reasoning for Neat.
@@ D.3 Commonsense Reasoning
We provide hyperparameters settings of Neat for commonsense reasoning task in Table 9. We follow the hyperparameters settings in MiLoRA [Wang et al., 2024c]. We limit all samples to a maximum of 256 tokens. For evaluation, we set a maximum token number of 32.
D.4 Arithmetic Reasoning
Table 10: Hyperparameter of arithmetic reasoning for Neat.
@@ D.4 Arithmetic Reasoning
We provide hyperparameters settings of Neat for arithmetic reasoning task in Table 10. We follow the hyper-parameters settings in MiLoRA [Wang et al., 2024c]. We limit all samples to a maximum of 2048 tokens. For evaluation, we set a maximum token number of 256 on GSM8K [Cobbe et al., 2021] dataset. On MATH [Hendrycks et al., 2021], we set the maximum new token to 512.
Appendix E Datasets
@@ E.1 Image Classification
E.2 Natural Language Understanding
@@
Figure 5: Implementation of introducing more depths to Neatt. We insert multiple intermediate layers into the layers from vanilla Neat, with non-linear activation in between. The depth is described as the number of layers in Neat, with vanilla Neat having a depth of 2 (i.e. the input and output layers).
@@ E.3 Commonsense Reasoning
Figure 6: An example of the actual implementation applying 4 layers in Neat (depth = 4) with Pytorch.
E.4 Arithmetic Reasoning
@@
Figure 7: Accuracy on the StanfordCars, FGVC and Oxford-Pets dataset with a wider range of depths for the neural network used in Neat. We choose depth equals to 2, 4, 6, 8 and 10 layers in the figure. In many cases, Neat with more depths outperforms vanilla Neat.
Appendix F Depths of Neat
We provide a comprehensive explanation of our approach to increasing the depth of the adaptation neural network in Neat. As depicted in Fig. 5, we introduce multiple deeply stacked intermediate layers between the layers of the vanilla Neat. These intermediate layers are essentially small adapters with a minimal parameter count ($\mathbb{R}^{r\times r}$, where $r$ is the hidden layer dimension), and we retain non-linear activations between them, as proposed by Neat. The adaptation process begins by feeding the weight matrix $\mathbf{W}^{0}$—the initialized value of the adaptation target $\mathbf{W}$—into Neat’s input layer. After undergoing multiple non-linear transformations through the intermediate layers, the final layer projects $\mathbf{W}^{0}$ back to its original shape, producing the adaptation result $\Delta W$. Throughout this process, the adaptation target remains fixed, while all the intermediate layers, as well as the input and output layers in Neat, are trainable parameters.
Furthermore, an implementation example of Neat with four layers using the PyTorch library is illustrated in Fig. 6. As previously mentioned, we apply non-linear activations (ReLU in this case) to model more complex transformations. The intermediate layers have the same shape, $\mathbb{R}^{r\times r}$, which adds minimal overhead compared to $\mathbf{A}\in\mathbb{R}^{d_{2}\times r}$ and $B\in\mathbb{R}^{r\times d_{2}}$—the input and output layers, respectively, which are also present in the vanilla Neat. Since $d_{2}$ is typically in the range of hundreds to thousands, while $r$ is commonly set to 8, 16, or 32, the parameter efficiency of Neat with deeper layers remains comparable to that of vanilla Neat without the intermediate layers. As shown, we first transform $\mathbf{W}^{0}$ into the desired adaptation result $\Delta W$ and subsequently use $\Delta W$ to perform the actual computation on the input data. The use of residuals is based on empirical observations, as incorporating residual connections in the adaptation process results in faster convergence, more stable loss curves, and significantly improved overall performance.
Meanwhile, to provide a comprehensive understanding of the impact of depth on model performance, we present the results of different depth configurations on the GLUE benchmark in Tab 15. Our analysis reveals that increasing the number of layers in the neural network architecture introduced by Neat consistently enhances performance while maintaining the same level of parameter efficiency. Furthermore, we extend our experiments to vision tasks, where even deeper networks are explored, and summarize the results in Fig. 7. We find that when the depth is increased to 8 and 10 layers, the accuracy slightly decreases compared to the 6-layer model but remains higher than that of the 2-layer configuration. A possible explanation is that as depth increases—particularly at 10 layers—the training process becomes more challenging, possibly requiring more careful hyper-parameter tuning. In general, these findings demonstrate that increasing the depth of the network is consistently beneficial for improving performance across both language and vision tasks.
Table 15: Accuracy comparison of Neat using RoBERTa-base with different depth configurations on the GLUE benchmark. The highest accuracy of methods per category are in bold. “AVG” means the average accuracy of all datasets.

### v3

@@ 7 Conclusion
In this work, we propose PEANuT, a novel parameter-efficient fine-tuning (PEFT) method that introduces nonlinear transformations to enhance model adaptation while maintaining efficiency. By incorporating a lightweight neural network that models cumulative weight updates as functions of the pre-trained weights, PEANuT effectively captures complex, nonlinear structures in the weight space, allowing for more expressive and accurate adaptation to downstream tasks. Our theoretical analysis supports the efficacy of PEANuT, demonstrating that it can achieve greater or equivalent expressiveness compared to existing LoRA, a popular and state-of-the-art PEFT method, with fewer number of parameters. Through extensive experiments on four benchmarks encompassing over twenty datasets with various pre-trained backbones, PEANuT demonstrated superior performance on both NLP and vision tasks compared to existing state-of-the-art methods.
Appendix A Appendix
Appendix B Details of Theoretical Results
@@
Take $\mathbf{A}^{\#}=\sigma(\mathbf{W}^{0}\bm{\Theta}_{1}^{*})\in\mathbb{R}^{d_{1}\times r}$ and $\mathbf{B}^{\#}=\bm{\Theta}_{2}^{*}\in\mathbb{R}^{r\times d_{2}}$, where $(\bm{\Theta}_{1}^{*},\bm{\Theta}_{2}^{*})=\operatorname*{arg\,min}_{\bm{\Theta}_{1}\in\mathbb{R}^{d_{2}\times r},\bm{\Theta}_{2}\in\mathbb{R}^{r\times d_{1}}}\mathcal{L}(\mathcal{D}_{\text{train}};\mathbf{W}^{0}+f(\mathbf{W}^{0};(\bm{\Theta}_{1},\bm{\Theta}_{2})))$. The conclusion follows from
@@
B.2 Theoretical Analysis of PEANuT under sinusoid activation function
Here we consider a sinusoid activation function $\sigma_{\textnormal{p}}(x)=\sin(2\pi x)$ (gashler2014training) and design $f(\mathbf{W}^{0};\bm{\theta})=\sigma_{\textnormal{p}}(\mathbf{W}^{0}\bm{\Theta}_{1})\bm{\Theta}_{2}$ with $\bm{\theta}=(\bm{\Theta}_{1},\bm{\Theta}_{2})$. With this periodic activation function, we can show a stronger result that PEANuT has expressivity (almost) greater than or equal to a LoRA with more parameters when $d_{1}\gg d_{2}$.
Proposition B.1 (Expressivity of PEANuT with Sine Activation).
@@ B.2 Theoretical Analysis of PEANuT under sinusoid activation function
Proposition B.1 shows that the class of updates $\Delta\mathbf{W}=\sigma_{\textnormal{p}}(\mathbf{W}^{0}\bm{\Theta}_{1})\bm{\Theta}_{2}$ by PEANuT with $2rd_{2}$ parameters is dense in the class of updates $\Delta\mathbf{W}=\mathbf{A}\mathbf{B}$ by LoRA with $r(d_{1}+d_{2})$ parameters. When $d_{2}\ll d_{1}$, this shows better parameter efficiency of PEANuT.
Examining the proof of Proposition B.1, it is straightforward to show that the result holds for any continuous and periodic activation function whose range contains an open interval centered at 0.
@@
This proof relies on Kronecker’s theorem (Theorem 7.9 in apostolmodular) from number theory, which shows that for all $j\in\mathbb{R}^{q}$, the fractional parts of $(ct_{1},ct_{2},\dots,ct_{q})^{\top}$ is dense in $0,1]^{q}$ over $c\in\mathbb{R}$, as long as $t_{1},\dots,t_{q}$ are linearly independent over the rationals.
@@
[…]
@@
Appendix C Hyperparameters
We provide the specific hyperparameters used in our experiments to ensure reproducibility. For most of our experiments, we use the standard implementation of PEANuT, which we refer to as vanilla PEANuT. The neural network architecture in vanilla PEANuT consists of only two layers: an input layer and an output layer. We select this approach because vanilla PEANuT offers the benefits of simplicity in implementation, a low parameter count, and sufficient adaptation power. Nonetheless, we dedicate Section 5.5 to exploring more complex adaptation networks and their effect on performance.
C.1 Image Classification
Hyperparameters for PEANuT for Fig. 5 are provided in Table 8. We tune the classification head and the backbone separately and provide detailed settings for each dataset. All weight decay values are not tuned and follow the settings from fourierft. The scaling factor $s$ is set to $1.0$. The hidden layer dimension $r$ for MHSA is set to 7 in the QV-setting, while both hidden layer dimensions for MHSA and MLP are set to 2 in the QV-MLP-setting described in Section 6.
Table 8: Hyperparameter of image classification for PEANuT.
@@ C.1 Image Classification
Learning Rate (PEANuT) 5E-3 1E-2 5E-3 1E-2 5E-3 1E-2 5E-3 5E-3
@@
Learning Rate (PEANuT) 5E-3 5E-3 5E-3 1E-2 5E-3 5E-3 1E-2 5E-3
@@
Table 9: Hyperparameter of commonsense reasoning for PEANuT.
@@
Table 10: Hyperparameter of arithmetic reasoning for PEANuT.
@@
C.2 Natural Language Understanding
We provide used hyper-parameters for PEANuT in natural language understanding on the GLUE benchmark in Table 11 and Table 12. The reported results are obtained when using a depth of 6 for PEANuT. The learning rates for the head and the backbone are tuned separately. The scaling factor $s$ is searched in $\{0.01,0.1,1.0\}$. For reproducibility, we fix the seed as 0. The hidden layer dimension $r$ is set to 8 in PEANuT-L and 1 in PEANuT-S. More specifically, we apply PEANuT to all layers in RoBERTa-base for PEANuT-L, while only applying PEANuT to layers $\{4,5,6,7,8,9,10,11\}$ for PEANuT-S to reduce the number of trainable parameters. The seed is fixed for reproducibility.
Table 11: Hyperparameter of GLUE benchmark for PEANuT-L.
Hyperparameter STS-B RTE MRPC CoLA SST-2 QNLI MNLI QQP
Optimizer AdamW
LR Schedule Linear
Learning Rate (PEANuT) 5E-3 5E-3 5E-3 1E-3 5E-3 1E-3 5E-3 5E-3
Learning Rate (Head) 5E-3 5E-3 5E-3 1E-3 5E-3 1E-3 5E-3 5E-3
Scaling 0.1 0.01 0.01 0.1 0.01 0.01 0.01 0.01
Max Seq. Len 512 512 512 512 512 512 512 512
Batch Size 64 32 64 64 32 32 32 64
Table 12: Hyperparameter of GLUE benchmark for PEANuT-S.
Hyperparameter STS-B RTE MRPC CoLA SST-2 QNLI MNLI QQP
Optimizer AdamW
LR Schedule Linear
Learning Rate (PEANuT) 5E-3 1E-3 5E-3 5E-3 5E-3 1E-3 5E-3 1E-3
Learning Rate (Head) 1E-3 1E-3 5E-3 1E-3 5E-3 1E-3 5E-3 1E-3
Scaling 0.1 1.0 0.01 0.1 0.01 0.1 0.01 1.0
Max Seq. Len 512 512 512 512 512 512 512 512
Batch Size 64 32 64 64 32 32 32 64
C.3 Commonsense Reasoning
We provide hyperparameters settings of PEANuT for commonsense reasoning task in Table 9. We follow the hyperparameters settings in MiLoRA (milora). We limit all samples to a maximum of 256 tokens. For evaluation, we set a maximum token number of 32.
C.4 Arithmetic Reasoning
We provide hyperparameters settings of PEANuT for arithmetic reasoning task in Table 10. We follow the hyper-parameters settings in MiLoRA (milora). We limit all samples to a maximum of 2048 tokens. For evaluation, we set a maximum token number of 256 on GSM8K (gsm8k) dataset. On MATH (MATH), we set the maximum new token to 512.
Appendix D Datasets
@@ D.1 Image Classification
D.2 Natural Language Understanding
@@ D.3 Commonsense Reasoning
D.4 Arithmetic Reasoning